# iConFormer: Dynamic Parameter-Efficient Tuning with Input-Conditioned Adaptation

### v1

@@ Introduction
The input-Conditioned Network (iCoN) is a key component of the iConFormer framework. Inspired by the concept of dynamic filter networks (Jia et al. 2016), iCoN employs a channel-wise mechanism to generate convolutional kernels that are tailored to the unique characteristics of each input. This approach enhances parameter efficiency while maintaining the flexibility needed to capture diverse features effectively. Formally, the iCoN module generates channel-wise convolutional kernels using input features from the MLP output, denoted as $\tilde{x}\in\mathbb{R}^{B\times M\times D}$ of (2). First, the feature $\tilde{x}$ is down-projected to $d$ channels and reshaped from $M$ into the spatial dimensions $\hat{H}\times\hat{W}$, which correspond to $H/P\times W/P$. This reshaping effectively rearranges the patches into a spatial grid to produce the feature $\hat{x}\in\mathbb{R}^{B\times\hat{H}\times\hat{W}\times d}$. Subsequently, the iCoN module generates dynamic convolutional kernels from the reshaped feature $\hat{x}$. This process is mathematically represented as:
@@
where $W_{dynamic}\in\mathbb{R}^{B\times D\times K\times K}$ is the dynamically generated kernel and $K$ is the kernel size. For our implementation, we set $K$ to 3 and $d$ to 64 ($d\ll D$). The dynamically generated convolution kernel $W_{dynamic}$ is then applied to feature $\hat{x}$, enabling a more precise capture of each input’s distinct features. Afterward, the feature is reshaped back to $\mathbb{R}^{B\times M\times d}$, followed by applying a non-linear activation function and up-projection to restore the channel dimension to the original size $D$. This process produces the final output of adapter as follows:
@@
By leveraging dynamically generated convolutional kernels, iConFormer refines the capture of input-specific features compared to conventional adapters, while maintaining parameter efficiency. This efficiency is achieved by employing a consistent set of parameters within the iCoN to generate convolutional kernels, even though the kernels are dynamically adapted for each specific input. Furthermore, the efficiency is enhanced by generating these dynamic kernels on a channel-wise basis, thereby minimizing the number of tunable parameters while preserving effective adaptation.
Consequently, iConFormer strikes a balance between effective feature extraction and computational efficiency relative to conventional PEFT methods, addressing their limitations and providing a robust solution for parameter-efficient adaptation.
Visual Analysis of Local Representation
@@
To evaluate the performance of capturing both local and global information, Fig. 3 visualizes the attention maps that offer insights on where the model focuses on. The attention maps are generated using attention rollout (Abnar and Zuidema 2020) with AdaptFormer (Chen et al. 2022) and iConFormer, employing Swin Transformer (Liu et al. 2021) as the backbone. Attention rollout computes token attentions by recursively multiplying attention matrices across layers, revealing how attention is distributed across different regions of the image. AdaptFormer adopts a standard pipeline of conducting down-projection, non-linear activation, and up-projection using static weight parameters that are not dynamically adjusted conditioned on input features. While the attention maps of AdaptFormer exhibit ambiguous and scattered attention distributions, with limitations in capturing fine-grained local features such as object edges, the attention maps of iConFormer are significantly more focused and well-aligned with objects.
@@
iConFormer utilizes Input-Conditioned Network (iCoN) to dynamically generate convolutional filters tailored to the input features. These convolutional kernels enable iConFormer to capture detailed spatial features while preserving overall contextual awareness. By focusing on these salient details, iConFormer demonstrates significant improvements in processing complex input data, leading to enhanced accuracy and robustness in dense prediction tasks.
Experiments
Experimental Settings
Datasets and Downstream Tasks.
To evaluate the effectiveness of iConFormer, we conduct comprehensive experiments on both image classification and dense prediction tasks, including monocular depth estimation, semantic segmentation, and instance segmentation. Implementation details can be found in the supplementary material. The datasets used in the experiments are as follows:
  • Image Classification: CIFAR-100 dataset (Krizhevsky, Hinton et al. 2009) consists of 50,000 training images and 10,000 validation images, each with a resolution of 32×32, categorized into 100 classes. The SVHN dataset (Netzer et al. 2011) includes over 600,000 labeled images for digit classification, comprising 73,257 training samples, 26,032 test samples, and 531,131 additional training images. The Food-101 dataset (Bossard, Guillaumin, and Van Gool 2014) contains 101,000 images across 101 food categories, with each category having 750 training samples and 250 test samples.
  • Monocular Depth Estimation: NYU-v2 (Silberman et al. 2012) with diverse indoor scenes and KITTI (Geiger, Lenz, and Urtasun 2012) with high-resolution outdoor driving scenes are benchmark datasets for depth estimation. For experiments, we used the standard splits and evaluate using Root Mean Square Error (RMSE). NYU-v2 images were cropped to 352 × 352 pixels, while KITTI images were cropped to 480 × 480 pixels.
  • Semantic Segmentation: ADE20K (Zhou et al. 2017) is a widely used semantic segmentation dataset with 20,000 training and 2,000 validation images. For our experiments, we utilized UperNet (Xiao et al. 2018) as the framework and evaluated performance using the mean Intersection over Union (mIoU) metric.
  • Instance Segmentation: MS COCO (Lin et al. 2014) is a prominent dataset for instance segmentation, with 118,000 training and 5,000 validation images. We used Cascade Mask R-CNN (Cai and Vasconcelos 2018; He et al. 2017) as a task-specific decoder and measured performance with Average Precision for bounding boxes ($\textrm{AP}_{Box}$) and masks ($\textrm{AP}_{Mask}$).
Pretrained Backbones.
For a fair comparison with FFT baseline and current PEFT methods, we used different pre-trained backbones depending on the tasks. In the semantic segmentation and instance segmentation tasks, Swin Transformer backbones (Liu et al. 2021), pre-trained on ImageNet-22k dataset (Deng et al. 2009), were used (Chen et al. 2019). Specifically, we used the Swin-Large backbone for semantic segmentation and the Swin-Base backbone for instance segmentation. For the monocular depth estimation, we utilized the standard Swin-V2-Base backbone (Liu et al. 2022) pre-trained using the MIM (Xie et al. 2022a). For the classification task, we adopted the ViT backbone (Dosovitskiy et al. 2021) pre-trained using MAE (He et al. 2022a).
@@
Table 1: Performance evaluation of image classification. We report the absolute Top-1 accuracy on the CIFAR-100, SVHN, and Food-101 datasets. indicates a learning rate reduced to 0.1×0.1$\times$0.1 × due to unstable training.
Method Params (M) CIFAR-100 SVHN Food-101
Full 86.04 (100%) 85.90 $97.67^{\dagger}$ $90.09^{\dagger}$
Linear 0.07 (0.08%) 69.83 66.91 69.74
VPT 0.08 (0.09%) 82.44 94.02 82.98
AdaptFormer 1.26 (1.46%) 85.90 96.89 87.61
Adapter 2.46 (2.86%) 86.65 97.09 86.89
iConFormer 1.71 (1.98%) 86.94 97.38 87.97
@@
Partial-l 12.62 (14.52%) 0.492 0.797 0.928 0.307 0.925 0.346
BitFit 0.14 (0.16%) 0.823 0.969 0.992 0.144 0.466 0.169
LN-tuning 0.05 (0.06%) 0.802 0.963 0.990 0.152 0.496 0.180
Adapter 3.11 (3.45%) 0.901 0.987 0.997 0.104 0.361 0.130
AdaptFormer 1.55 (1.76%) 0.845 0.975 0.994 0.134 0.432 0.159
LoRA 3.08 (3.42%) 0.439 0.733 0.885 0.402 1.045 0.395
@@
Baseline methods.
For the image classification task, we used the same set of comparison models as in (Chen et al. 2022), and additionally included the Adapter method from (Houlsby et al. 2019). In the monocular depth estimation, we included comparisons with partial tuning such as BiTFiT (Zaken, Ravfogel, and Goldberg 2021), LN-Tuning (Qi et al. 2022). We also evaluated against Partial-l (Yosinski et al. 2014), which fine-tunes only the final block of the backbone and parameters outside the backbone. For adapter-based methods, we compared with recent approaches, including Adapter (Houlsby et al. 2019), AdaptFormer (Chen et al. 2022), LoRA (Hu et al. 2021), and LoRand (Yin et al. 2023). In the semantic segmentation and instance segmentation tasks, we configured the Adapter model under the same settings as described in (Yin et al. 2023) to ensure a fair comparison. Across all tasks, we also included full fine-tuning (FFT) for performance evaluation.
Main Results
@@
Adapter-B 32.04 (16.13%) 46.23 %
Adapter-T 16.04 (8.08%) 43.51 %
AdaptFormer 2.34 (1.18%) 49.85 %
LoRA 4.57 (2.31%) 49.48 %
LoRand 3.59 (1.84%) 50.67 %
iConFormer 3.26 (1.65%) 50.82%
@@
Adapter-B 14.38 (16.13%) 46.50 % 40.20 %
Adapter-T 7.20 (8.08%) 43.20 % 38.70 %
AdaptFormer 1.60 (1.79%) 48.90 % 42.50 %
LoRA 3.08 (3.43%) 47.50 % 41.50 %
LoRand 2.39 (2.76%) 51.10 % 44.10 %
iConFormer 2.48 (2.78%) 52.70% 45.70%
Image classification.
@@
We evaluate various fine-tuning approaches using backbones pre-trained via self-supervised learning paradigms (Grill et al. 2020; Chen and He 2021; Choi et al. 2023b, a), as detailed in Table 1. The results demonstrate that iConFormer consistently outperforms linear probing, Visual Prompt Tuning (VPT) methods, and recently proposed adapter-based techniques. Specifically, iConFormer achieves performance improvements of 4.5%, 3.36%, and 4.99% over VPT on the image benchmarks CIFAR-100, SVHN, and Food-101, respectively. Furthermore, when compared to recent adapter-based methods such as Adapter and AdaptFormer, iConFormer shows up to 1.04%, 0.49%, and 1.08% higher accuracy, respectively. Notably, iConFormer also surpasses the full fine-tuning approach by more than 1% Top-1 accuracy on the CIFAR-100 dataset. In summary, iConFormer is not only highly efficient compared to other adapter-based methods but also achieves superior performance with up to 2% of the parameters used in full fine-tuning.
@@
Table 5: Ablation study of input-Conditioned kernel size in iCoN. We perform a quantitative comparison of different kernel sizes on the KITTI dataset. Results with the symbol $\uparrow$ / $\downarrow$ indicate the higher/lower is the better.
@@
$3\times 3$ 2.48 (2.68%) 0.964 0.996 0.999 0.061 2.302 0.089
$5\times 5$ 4.07 (4.48%) 0.964 0.996 0.999 0.060 2.314 0.090
$7\times 7$ 6.47 (6.93%) 0.963 0.996 0.999 0.060 2.320 0.090
Monocular Depth Estimation.
Table 2 present the performance results for the NYU-v2 datasets. As shown in the tables, iConFormer outperforms other parameter efficient fine-tuning methods in all metrics, with the RMSE value being within 0.04 of the full fine-tuning performance. Moreover, iConFormer shows an RMSE improvement of up to 0.2 compared to partial tuning methods, and an enhancement of up to 0.3 RMSE compared to adapter-based methods such as Adapter, AdaptFormer, and LoRA. These results suggest that iConFormer’s capability to generate and apply input-conditioned kernels significantly contributes to its precision in depth estimation tasks. The performance evaluation on the KITTI dataset is presented in the supplemental material.
Semantic Segmentation.
We present the results of the semantic segmentation task on the ADE20K dataset in Table 3. By fine-tuning fewer than 3.3 million backbone parameters, iConFormer achieves 50.82% mIoU on ADE20K, which is approximately 0.7% lower than Full fine-tuning. Moreover, iConFormer requires fewer parameters for tuning compared to previously proposed adapter-based methods while still achieving superior performance. These results suggest that iConFormer effectively utilizes a limited subset of parameters to capture task-specific information and learn detailed features.
Instance Segmentation.
Table 4 presents the instance segmentation results on the COCO dataset. iConFormer demonstrates significant performance gains by training only 2.78% of the total backbone parameters, surpassing both existing adapter-based methods and full fine-tuning. Specifically, it achieves a 0.8% improvement in $\text{AP}_{box}$ and a 0.7% improvement in $\text{AP}_{mask}$ compared to full fine-tuning. These results effectively reveal the advantages of the proposed method and demonstrate its superiority over full fine-tuning in terms of both storage efficiency and performance. Additionally, these findings suggest that iConFormer optimizes resource utilization through its dynamic kernel approach.
Ablation Studies
We conducted ablation studies to explore various aspects of iConFormer and identify the key factors that contribute to its performance. All ablation experiments were performed using the dense tasks.
input-Conditioned Kernel Size.
In Table 5, we conducted an ablation study on the KITTI dataset by varying the size of the input-conditioned kernel. We experimented with kernel sizes of 3$\times$3, 5$\times$5, and 7$\times$7. The results demonstrate that iConFormer performs well with a 3$\times$3 kernel across most evaluation metrics. Notably, the RMSE slightly improves as the kernel size decreases, with 3$\times$3 kernel achieving the lowest RMSE. This observation indicates that 3$\times$3 kernel captures essential local features effectively while maintaining computational efficiency. Given that the performance across kernel sizes is quite similar, 3$\times$3 kernel was adopted for its efficiency, providing a balanced trade-off between accuracy and computational cost for the input-conditioned kernels of iCoN.
Table 6: Ablation study of the sequential and parallel configurations on dense prediction tasks. Results are presented in the order of NYU-v2, ADE20K, and COCO datasets, from top to bottom.
@@
Parallel 2.48 (2.68%) 0.357
Sequential 2.48 (2.68%) 0.342
[…]
Parallel 3.26 (1.65%) 50.25 %
Sequential 3.26 (1.65%) 50.82 %
Configuration Params (M) $\textrm{AP}_{Box}\uparrow$ $\textrm{AP}_{Mask}\uparrow$
Parallel 2.48 (2.78%) 51.0 % 44.4 %
Sequential 2.48 (2.78%) 52.7% 45.7%
iConFormer Configuration.
We investigated the iConFormer configuration by comparing its sequential and parallel instances, as illustrated in Figure 4, where the distinction is based on the placement of the Adapter within the Transformer block. As demonstrated in Table 6, the sequential form of iConFormer significantly outperforms the parallel form across all dense tasks. The reason might be: (1) the sequential design processes each layer’s output in a progressive manner, facilitating deeper feature representations and gradual refinement of complex patterns; (2) the parallel design processes outputs simultaneously, which results in limited inter-layer interaction, weakening the information flow and hindering the model’s capacity to capture intricate features. Therefore, we adopted the sequential design as the default configuration for iConFormer, given its demonstrated superior performance.
Conclusion

### v2

@@ 4.2 Input-Conditioned Network (iCoN)
The iCoN is a key component of the iConFormer framework. Inspired by the concept of dynamic filter networks [36], iCoN employs a channel-wise mechanism to dynamically generate convolutional kernels that are tailored to the unique characteristics of each input. This approach enhances parameter efficiency while maintaining the flexibility needed to capture diverse features effectively. Formally, the iCoN module generates channel-wise convolutional kernels using input features from the MLP output, denoted as $\tilde{x}\in\mathbb{R}^{B\times M\times D}$ of (2). First, the feature $\tilde{x}$ is down-projected to $d$ channels and reshaped from $M$ into the spatial dimensions $\hat{H}\times\hat{W}$, which correspond to $H/P\times W/P$. This reshaping rearranges the patches into a spatial grid, producing the feature $\hat{x}\in\mathbb{R}^{B\times\hat{H}\times\hat{W}\times d}$. Subsequently, the iCoN module generates dynamic convolutional kernels from the reshaped feature $\hat{x}$. This process is mathematically represented as:
@@
where $W_{dynamic}\in\mathbb{R}^{B\times d\times K\times K}$ is the dynamically generated kernel and $K$ is the kernel size. For our implementation, we set $K$ to 3 and $d$ to 64 ($d\ll D$).
Figure 4 illustrates the process of dynamically generating convolutional kernels in the iCoN module. It first applies spatial average pooling to extract global contextual information from $\hat{x}$. The pooled features are then passed through a lightweight transformation, which learns a mapping function to parameterize the convolutional kernel weights. Finally, the transformed features are reshaped into the channel-wise convolutional kernel $W_{dynamic}$. The convolution kernel generated conditioned on the input feature $\hat{x}$ is then applied to $\hat{x}$ through a channel-wise convolution operation. Afterward, the feature is reshaped back to $\mathbb{R}^{B\times M\times d}$, followed by applying a non-linear activation function and up-projection to restore the channel dimension to the original size $D$. This process produces the final output of the adapter as follows:
@@
By dynamically generating convolutional kernels, iCoN allows the model to effectively adapt to the input structure while enriching the feature representation with both global context and local details. Here, it should be noted that the standard convolution operation can also be applied to the Transformer encoder to inject the locality, but the dynamic adaptation to the input instance is more crucial to the PEFT framework for various downstream tasks. In contrast to standard convolutional kernels, iCoN dynamically adjusts its kernels based on the spatial structure of each input, resulting in more precise and context-aware feature extraction. This mechanism allows iCoN to capture high-frequency variations and subtle patterns that the standard convolution layers struggle to represent effectively. This is also validated in experiments (Table 7).
It is worthy of noting that while our approach generates dynamic kernels, it remains computationally comparable to conventional adapter-based PEFT methods [10, 33]. This efficiency is primarily attributed to two factors. First, while iCoN generates dynamic kernels, the network responsible for the kernel generation retains static learnable parameters, ensuring that only the output kernels adapt to instance without significantly increasing parameter overhead. Second, to further reduce computational complexity while preserving adaptive capacity, iCoN employs the channel-wise convolutions instead of the original convolutions, as shown in Figure 4.
Figure 5: Comparison of Attention Maps from AdaptFormer and iConFormer. We visualize the attention maps using attention rollout [1]. The top row represents input images, and the middle and bottom rows present the attention maps generated by AdaptFormer [10] and iConFormer, respectively, with both using Swin Transformer backbone [47]. iConFormer more accurately delineates object regions and captures fine-grained semantics, compared to the AdaptFormer.
4.3 Visual Analysis of Local Representation
@@
To evaluate the effectiveness of capturing both local and global information, Figure 5 visualizes the attention maps that provide insight where the model focuses. The attention maps are generated using attention rollout [1] with AdaptFormer [10] and iConFormer, employing the Swin Transformer [47] as the backbone. Attention rollout computes token attentions by recursively multiplying attention matrices across layers, revealing how attention is distributed across different regions of an input image. AdaptFormer adopts a standard pipeline consisting of down-projection, non-linear activation, and up-projection based on static weight parameters that are not dynamically adjusted conditioned on input features. While the attention maps of AdaptFormer exhibit ambiguous and scattered attention distributions with limitations in precisely capturing local features such as object edges, the attention maps of iConFormer are significantly more focused and better aligned with objects.
@@
The iCoN used in the iConFormer dynamically generates convolutional kernels tailored to the input features, enabling the iConFormer to capture detailed spatial features while preserving overall contextual awareness. By focusing on these salient details, the proposed method demonstrates significant improvements in processing complex input data, leading to enhanced accuracy and robustness in dense prediction tasks.
5 Experiments
@@
Table 1: Performance evaluation of image classification. We report the absolute Top-1 accuracy on the CIFAR-100, SVHN, and Food-101 datasets. indicates a learning rate reduced to 0.1×0.1$\times$0.1 × due to unstable training. Additionally, FLOPs are measured with a batch size of 2.
Method Params (M) FLOPs (G) CIFAR-100 SVHN Food-101
Full 86.04 (100%) 35.16 85.90 97.67 90.09
Linear 0.07 (0.08%) 35.16 69.83 66.91 69.74
VPT [35] 0.08 (0.09%) 35.35 82.44 94.02 82.98
AdaptFormer [10] 1.26 (1.46%) 35.63 85.90 96.89 87.61
Adapter [33] 2.46 (2.86%) 36.13 86.65 97.09 86.89
iConFormer 1.71 (1.98%) 35.65 86.94 97.38 87.97
5.1 Experimental Settings
@@
Partial-l [66] 12.62 (14.52%) 0.492 0.797 0.928 0.307 0.925 0.346
BitFit [67] 0.14 (0.16%) 0.823 0.969 0.992 0.144 0.466 0.169
LN-tuning [51] 0.05 (0.06%) 0.802 0.963 0.990 0.152 0.496 0.180
Adapter [33] 3.11 (3.45%) 0.901 0.987 0.997 0.104 0.361 0.130
AdaptFormer [10] 1.55 (1.76%) 0.845 0.975 0.994 0.134 0.432 0.159
LoRA [34] 3.08 (3.42%) 0.439 0.733 0.885 0.402 1.045 0.395
@@
Adapter-B [33] 32.04 (16.13%) 46.23 %
Adapter-T [33] 16.04 (8.08%) 43.51 %
AdaptFormer [10] 2.34 (1.18%) 49.85 %
LoRA [34] 4.57 (2.31%) 49.48 %
LoRand [65] 3.59 (1.84%) 50.67 %
iConFormer 3.26 (1.65%) 51.17%
@@
Adapter-B [33] 14.38 (16.13%) 46.50 % 40.20 %
Adapter-T [33] 7.20 (8.08%) 43.20 % 38.70 %
AdaptFormer [10] 1.60 (1.79%) 48.90 % 42.50 %
LoRA [34] 3.08 (3.43%) 47.50 % 41.50 %
LoRand [65] 2.39 (2.76%) 51.10 % 44.10 %
iConFormer 2.48 (2.78%) 52.90% 45.90%
Datasets and Downstream Tasks
To evaluate the performance of iConFormer, we conducted comprehensive experiments on both image classification and dense prediction tasks, including monocular depth estimation, semantic segmentation, and instance segmentation. Implementation details can be found in the supplementary material. The datasets used in the experiments are as follows:
  • Image Classification: CIFAR-100 dataset [38] consists of 50,000 training images and 10,000 validation images, each with a resolution of 32×32, categorized into 100 classes. The SVHN dataset [50] includes over 600,000 labeled images for digit classification, comprising 73,257 training samples, 26,032 test samples, and 531,131 additional training images. The Food-101 dataset [4] contains 101,000 images across 101 food categories, with each category having 750 training samples and 250 test samples.
  • Monocular Depth Estimation: NYU-v2 [55] with diverse indoor scenes and KITTI [23] with high-resolution outdoor driving scenes are benchmark datasets for depth estimation. For experiments, we used the standard splits and evaluate using Root Mean Square Error (RMSE). NYU-v2 images were cropped to 352 × 352 pixels, while KITTI images were cropped to 480 × 480 pixels.
  • Semantic Segmentation: ADE20K [68] is a widely used semantic segmentation dataset with 20,000 training and 2,000 validation images. For our experiments, we utilized UperNet [59] as the framework and evaluated performance using the mean Intersection over Union (mIoU) metric.
  • Instance Segmentation: MS COCO [45] is a prominent dataset for instance segmentation, with 118,000 training and 5,000 validation images. We used Cascade Mask R-CNN [5, 28] as a task-specific decoder and measured performance with Average Precision for bounding boxes ($\textrm{AP}_{Box}$) and masks ($\textrm{AP}_{Mask}$).
Pretrained Backbones
For a fair comparison with FFT baseline and current PEFT methods, we used different pre-trained backbones depending on the tasks. In the semantic segmentation and instance segmentation tasks, Swin Transformer backbones [47], pre-trained on ImageNet-22k dataset [19], were used [8]. Specifically, we used the Swin-Large backbone for semantic segmentation and the Swin-Base backbone for instance segmentation. For the monocular depth estimation, we utilized the standard Swin-V2-Base backbone [48] pre-trained using the MIM [61]. For the classification task, we adopted the ViT backbone [22] pre-trained using MAE [30].
Baseline Methods
For the image classification task, we used the same set of comparison models as [10], and additionally included the Adapter method from [33]. In the monocular depth estimation, we included comparisons with partial tuning methods such as BiTFiT [67], LN-Tuning [51]. We also evaluated against Partial-l [66], which fine-tunes only the final block of the backbone and parameters outside the backbone. For comparison with adapter-based methods, we included recent approaches such as Adapter [33], AdaptFormer [10], LoRA [34], and LoRand [65]. In the semantic segmentation and instance segmentation tasks, we configured the Adapter [33] following [65] for a fair comparison, setting the intermediate layer dimension to half of the input dimension for ‘Adapter-B’ and to a quarter for ‘Adapter-T’. Additionally, we included ‘Fixed’ in all dense prediction tasks, freezing the pre-trained Transformer encoder while training other parts of the architecture (i.e., task decoder). Across all tasks, we also included ‘Full’, which indicates full fine-tuning (FFT) as an upper bound on performance.
5.2 Main Results
Image Classification
@@ 5.2 Main Results
We evaluated various fine-tuning approaches using ViT backbone [22] pre-trained via self-supervised learning paradigms [26, 11, 16, 15], as detailed in Table 1. The results demonstrate that the iConFormer consistently outperforms linear probing, Visual Prompt Tuning (VPT) methods, and recently proposed adapter-based techniques. Specifically, the iConFormer achieves performance improvements of 4.5%, 3.36%, and 4.99% over VPT on the image benchmarks CIFAR-100, SVHN, and Food-101, respectively. Furthermore, when compared to recent adapter-based methods such as Adapter [33] and AdaptFormer [10], iConFormer shows up to 1.04%, 0.49%, and 1.08% higher accuracy, respectively. Notably, iConFormer also surpasses the FFT approach by more than 1% Top-1 accuracy on the CIFAR-100 dataset. Additionally, while consistently delivering better performance, iConFormer demonstrates computational efficiency with 35.65 GFLOPs, which is comparable to AdaptFormer (35.63G) and lower than Adapter (36.13G). In summary, our approach outperforms recent adapter-based approaches with similar computational efficiency, and provides comparable performance to the FFT despite tuning only 2% of the parameters used in the FFT.
Monocular Depth Estimation
Table 2 presents the performance results for the NYU-v2 datasets. As shown in the tables, the iConFormer outperforms other PEFT methods in all metrics, with the RMSE value being within 0.04 of the FFT performance. Moreover, the iConFormer shows an RMSE improvement of up to 0.2 compared to partial tuning methods, and an enhancement of up to 0.3 RMSE compared to adapter-based methods such as Adapter [33], AdaptFormer [10], and LoRA [34]. These results suggest that iConFormer’s capability to generate and apply input-conditioned kernels significantly contributes to the performance in the monocular depth estimation task. Additional results on the KITTI dataset are presented in the supplemental material.
Semantic Segmentation
We present the results of the semantic segmentation task on the ADE20K dataset [68] in Table 3. By fine-tuning fewer than 3.3 million backbone parameters, the proposed method achieves 51.17% mIoU on ADE20K, which is about 0.3% lower than the FFT. Moreover, the iConFormer requires fewer learnable parameters compared to most adapter-based methods while still achieving superior performance. These results suggest that iConFormer effectively utilizes a limited subset of parameters to capture task-specific information and learn detailed features.
Instance Segmentation
Table 4 presents the instance segmentation results on the COCO dataset. Our method demonstrates significant performance gains by training only 2.78% of the total backbone parameters, surpassing both existing adapter-based methods and FFT. Specifically, it achieves 1.0% improvement in $\text{AP}_{box}$ and 0.9% improvement in $\text{AP}_{mask}$ compared to the FFT. These results reveal the advantages of the proposed method and demonstrate its superiority over the FFT in terms of both storage efficiency and performance. Additionally, these findings suggest that iConFormer optimizes resource utilization through its dynamic kernel approach.
Table 5: Ablation study of the sequential and parallel configurations on dense prediction tasks. Results are presented in the order of NYU-v2, ADE20K, and COCO, from left to right.
Configuration RMSE ↓ mIoU ↑ $\textrm{AP}_{Box}\uparrow$ $\textrm{AP}_{Mask}\uparrow$
Parallel 0.357 50.85 % 51.20 % 44.60 %
Sequential 0.342 51.17 % 52.90 % 45.90 %
@@
Table 6: Ablation study of input-Conditioned kernel size in iCoN. We perform a quantitative comparison of different kernel sizes across dense prediction datasets. Results with the symbol $\uparrow$ / $\downarrow$ indicate the higher/lower is the better.
Kernel Size KITTI ADE20K COCO
Params (M) RMSE ↓ Params (M) mIoU ↑ Params (M) $\textrm{AP}_{Box}\uparrow$ $\textrm{AP}_{Mask}\uparrow$
3 $\times$ 3 2.48 (2.68%) 2.302 3.26 (1.65%) 51.17 % 2.48 (2.78%) 52.90 % 45.90 %
5 $\times$ 5 4.07 (4.48%) 2.314 4.86 (2.43%) 50.95 % 4.07 (4.49%) 52.60 % 45.70 %
7 $\times$ 7 6.47 (6.93%) 2.320 7.26 (3.59%) 50.87 % 6.47 (6.94%) 52.60 % 45.70 %
Table 7: Ablation study on the effect of the Input-Conditioned Kernel in dense prediction tasks. Results are presented in the order of NYU-v2, ADE20K, and COCO datasets, from top to bottom. Both kernel types uses a 3$\times$3 kernel size for all experiments.
@@
     Standard Conv      2.44 (2.64%)      1.029
     input-Conditioned Conv      2.48 (2.68%)      0.342
@@
     Standard Conv      3.25 (1.64%)      50.02 %
     input-Conditioned Conv      3.26 (1.65%)      51.17 %
Kernel Type Params (M) $\textrm{AP}_{Box}\uparrow$ $\textrm{AP}_{Mask}\uparrow$
Standard Conv 2.46 (2.76%) 50.20 % 43.50 %
input-Conditioned Conv 2.48 (2.78%) 52.90 % 45.90 %
5.3 Ablation Studies
We conducted ablation studies to explore various aspects of the iConFormer and identify key factors that contribute to its performance. All ablation experiments were conducted using the dense predictions tasks.
iConFormer Configuration
We investigated the performance by comparing sequential and parallel configurations, as illustrated in Figure 3, where the distinction is based on the placement of the Adapter within the Transformer block. As demonstrated in Table 5, the sequential form significantly outperforms the parallel form for all dense tasks. The reason might be: (1) the sequential design processes each layer’s output in a progressive manner, facilitating deeper feature representations and gradual refinement of complex patterns; (2) the parallel design processes outputs simultaneously, which results in limited inter-layer interaction, weakening the information flow and hindering the model’s capacity to capture intricate features. Therefore, we adopted the sequential design as the default configuration for iConFormer, given its demonstrated superior performance.
input-Conditioned Kernel Size
In Table 6, we present an ablation study on the size of the input-conditioned convolution on dense prediction tasks. Experiments with kernel sizes of 3$\times$3, 5$\times$5, and 7$\times$7 demonstrate that the 3×3 kernel consistently achieves competitive results with a relatively small number of parameters. Notably, on the KITTI dataset, the RMSE slightly improves as the kernel size decreases, with 3$\times$3 kernel achieving the lowest RMSE. Similarly, on the ADE20K and COCO datasets, the 3$\times$3 kernel consistently outperforms the 5$\times$5 and 7$\times$7 variants in both mIoU and AP. These results indicate that 3$\times$3 kernel captures essential local features effectively while maintaining computational efficiency. Given that the performance across kernel sizes is quite similar, 3$\times$3 kernel was adopted for its efficiency, providing a balanced trade-off between accuracy and computational cost for the input-conditioned kernels of iCoN.
Effect of input-Conditioned Kernel
We investigated the effect of using the input-conditioned convolution $W_{dynamic}$ of (6) in the iCoN for dense prediction tasks. In Table 7, we compared the performance when using the standard convolution and the input-conditioned convolution for the NYU-v2 dataset (monocular depth estimation), the ADE20K dataset (semantic segmentation), and the COCO dataset (instance segmentation). For a fair comparison, we set to $3\times 3$ kernel for both cases, ensuring the same local receptive field. The input-conditioned kernel consistently outperforms the standard variant for all tasks, improving mIoU by about 1.2% on ADE20K and AP by about 2.5% on COCO, and reducing RMSE by about 0.7 on NYU-v2.
To further analyze the performance of the standard convolution, we extended the comparison to existing PEFT methods. In Table 2, most existing PEFT methods achieve a lower RMSE than the standard convolution. In Table 3 and 4, compared to LoRand [65], the standard convolution achieves approximately 0.6% lower mIoU on ADE20K and about 1% lower AP on COCO, respectively. This indicates that simply applying the standard convolution to our framework is not so effective and the adaptive nature of the input-conditioned convolution, where kernel weights are dynamically modulated for input features, is more crucial to capture local details in dense prediction tasks.
6 Conclusion